OpenStudy (meilendurcer):

Which statement is an example of the Identity Property of Multiplication? 8 • 0 = 0 8 • 1 = 8 8 • –1 = –8 –8 • –1 = 8

OpenStudy (anonymous):

The identity property of multiplication, also called the multiplication property of one says that a number does not change when that number is multiplied by 1. So, which one do you think it is?

OpenStudy (meilendurcer):

i think it's C

OpenStudy (anonymous):

No, read it again please. 1, not -1...

OpenStudy (meilendurcer):

My first choice was A

OpenStudy (anonymous):

That's also wrong, it's specifically about n*1=n. Which choice reflects that?

OpenStudy (meilendurcer):

D?

OpenStudy (anonymous):

No, it's B... Please read it again, N*1=N, so 8*1=8.

OpenStudy (meilendurcer):

ahhh now i get it thanks

OpenStudy (lopus):

Commutative property: When two numbers are multiplied together, the product is the same regardless of the order of the multiplicands. For example 4 * 2 = 2 * 4 Associative Property: When three or more numbers are multiplied, the product is the same regardless of the grouping of the factors. For example (2 * 3) * 4 = 2 * (3 * 4) Multiplicative Identity Property: The product of any number and one is that number. For example 5 * 1 = 5. \[\checkmark\] Distributive property: The sum of two numbers times a third number is equal to the sum of each addend times the third number. For example 4 * (6 + 3) = 4*6 + 4*3

Directrix (directrix):

Use the attached chart of properties to review. Hope it helps.